OpenStudy (anonymous):

(square root of 3)(square root of 6)

OpenStudy (saifoo.khan):

Satellliteeeeeeeeeeee!

OpenStudy (anonymous):

\[\sqrt{3}\times \sqrt{6}=\sqrt{3\times 6}=\sqrt{18}=\sqrt{9\times 2}=\sqrt{9}\times \sqrt{2}=3\sqrt{2}\]

OpenStudy (saifoo.khan):

RIGHT^

OpenStudy (anonymous):

\[\color{red}{\text{saifooooooooooooooo!}}\]

OpenStudy (saifoo.khan):

xD